# Computing Wasserstein Barycenters through Gradient Flows

### v2

@@ C.2 Wasserstein Gradient of Functionals
Which is the update rule in [cuturi2014fast, Algorithm 2] with $\theta=\nicefrac{{2\alpha}}{{n}}$. Furthermore, we can also instantiate the fixed-point algorithm of [montesuma2023multi], for $\alpha=n/2$. We now highlight two differences of our work with respect these two strategies. First, our setting is more general. In fact, our algorithm is capable of handling ground metrics other than the squared Euclidean metric. Furthermore, our framework couples the barycenter objective with other functionals. Second, while [cuturi2014fast] assume that all samples of $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ are available at once, our algorithm assumes that these measures are available through sampling. This treatment is oftentimes called continuous in the literature [korotin2021continuous, korotin2022wasserstein], however, one still relies on samples from these measures rather than some truly continuous form.
C.3 Domain Adaptation
In this section, we offer a brief introduction to domain adaptation in classification problems which are the main application we explore in the main paper.
@@ C.3 Domain Adaptation
Statistical learning theory [vapnik1999overview] deals with the problem of learning a function $h:\mathcal{X}\rightarrow\mathcal{Y}$, such that it minimizes the risk of making a mistake, that is,
@@
There is a fundamental assumption at play in the theory of generalization under the ERM framework. Indeed, both $\mathcal{R}_{P}$ and $\hat{\mathcal{R}}_{P}$ evaluate the risk under a fixed probability measure $P$. However, machine learning models are oftentimes confronted with data following different, but related probability measures [quinonero2022dataset], a process known as distribution shift.
@@
Source-Only Training. A straightforward baseline in domain adaptation is the source-only training. Let $h=g\circ f$ be the composition of a feature extractor $g:\mathcal{X}\rightarrow\mathcal{U}$ and a feature classifier $f:\mathcal{U}\rightarrow\mathcal{Y}$. We can train $h$ end-to-end by minimizing,
@@
This training produces an embedding space $\mathcal{U}$ that is discriminative of the classes in the domain adaptation problem. Following previous work [montesuma2021wasserstein, montesuma2023multi, montesuma2024distillation], we perform adaptation in this space.
Wasserstein Barycenters and Domain Adaptation. The motivation for using Wasserstein barycenters in multi-source domain adaptation comes from a theoretical result of [redko2019optimal]. Indeed, by [redko2019optimal, Theorem 2],
@@
Plugging this back into inequality 42,
@@
4:  Compute $P=\text{Bar}(\lambda,\mathcal{Q})$
5:  Compute $\gamma^{\star}$ between barycenter and target using eq. 43
6:  Transport $(x_{i}^{(P)},y_{i}^{(P)})$ Using eq. 44
@@
Wasserstein Barycenter Transport $(p=2)$. [montesuma2021wasserstein] uses a two-step procedure for finding $P$. First, one solves,

### v3

@@ C.2 Wasserstein Gradient of Functionals
Which is the update rule in [cuturi2014fast, Algorithm 2] with $\theta=\nicefrac{{2\alpha}}{{n}}$. Furthermore, we can also instantiate the fixed-point algorithm of montesuma2023multi, for $\alpha=n/2$. We now highlight two differences of our work with respect these two strategies. First, our setting is more general. In fact, our algorithm is capable of handling ground metrics other than the squared Euclidean metric. Furthermore, our framework couples the barycenter objective with other functionals. Second, while cuturi2014fast assume that all samples of $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ are available at once, our algorithm assumes that these measures are available through sampling. This treatment is oftentimes called continuous in the literature korotin2021continuous, korotin2022wasserstein, however, one still relies on samples from these measures rather than some truly continuous form.
Appendix D Proof of Theorem 3.1
  1. 1.
    The exponential convergence given by the PL inequality,
    $\lVert\nabla_{W}\mathbb{B}(P)\rVert_{L_{2}(P)}^{2}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (42)
  2. 2.
    An uniform bound on $P\mapsto\mathbb{B}(P|\mathcal{Q})$ controlling the error of approximating $Q_{k}$ with $\hat{Q}_{k}$.
These results are available, respectively, in chewi2024statistical and portales2025sample. For completeness, we re-state and demonstrate them. We start with a modified version of [chewi2024statistical, Corollary 5.17],
Proposition D.1.
(Measure theoretic Gronwall’s Lemma chewi2024statistical) Let $\mathbb{F}:\mathcal{P}_{2}(\Omega)\rightarrow\mathbb{R}$ be a functional that satisfies the PL inequality (c.f. 42) with constant $C_{\text{PL}}>0$, then,
Proof.
Define $\phi(t)=\mathbb{F}(P_{t})-\mathbb{F}^{\star}$. Then,
$\partial_{t}\phi(t)$ $=\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla_{W}\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$
From Grönwall’s Lemma [chewi2024statistical, Lemma 5.16],
We now present the uniform approximation bound of portales2025sample for empirical barycenters,
Theorem D.1.
(Barycenter Empirical Approximation error portales2025sample) Let $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ be a finite family of measures on $\mathcal{P}_{2}(\mathcal{B}(0,R))$ for $R>0$. Approximate each $Q_{k}$ with $m$ i.i.d. samples $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$. Let $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{z_{i}^{(P)}}$. Let $p\in[1,+\infty)$ and $\epsilon\geq 0$. Then,
$\mathbb{E}\biggr[\underset{z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega}{\text{sup}}\,|\mathbb{B}_{\epsilon}(\hat{P}|\mathcal{Q})-\widehat{\mathbb{B}}_{\epsilon}(\hat{P}|\mathcal{Q})|\biggr]\leq C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$ (44)
where,
$C_{p,R}$ $=\begin{cases}8\sqrt{2}\int_{0}^{3(2R)^{p}}\sqrt{\log\biggr(2+\dfrac{32p(2R)^{p}}{s}\biggr)}ds&\epsilon=0,\\ 8\sqrt{2}\int_{0}^{(4p+1)(2R)^{p}}\sqrt{\log(2+\frac{64p(2R)^{p}}{s})}ds&\epsilon>0,\end{cases},$
$C_{d,n}$ $=n(d+1).$
With these bounds established, we can then prove our Theorem 3.1.
Proof.
Let $\tau>0$ be the current optimization step. We have the following,
[…]
From Equation 45 we can bound the expected empirical optimization gap,
$\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}|]\leq$ $\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\mathbb{B}(\hat{P}_{\tau})|]+\mathbb{E}[|\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}|]+\mathbb{E}[|\mathbb{B}^{\star}-\hat{\mathbb{B}}^{\star}|].$ (46)
Now, we can bound the first and third terms using the approximation error of portales2025sample (i.e., Equation 44),
$\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\mathbb{B}(\hat{P}_{\tau})|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$ (47)
$\mathbb{E}[|\hat{\mathbb{B}}^{\star}-\mathbb{B}^{\star}|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$
here, we recall that $\mathbb{B}^{\star}=\mathbb{B}(P^{\star})=\text{inf}_{P}\mathbb{B}(P)$. Returning to the second term, note that the true optimization gap $\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})-\mathbb{B}^{\star}$ is deterministic, as it does not involve sampling from the measures in $\mathcal{Q}$. This means that,
$\mathbb{E}[|\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}|]=|\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}|\leq e^{-C_{\text{PL}}\tau}(\hat{\mathbb{B}}(\hat{P}_{0})-\hat{\mathbb{B}}^{\star}).$ (48)
Combining Equations 46, 47 and 48 we get,
$\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}|]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}(\hat{P}_{0})-\mathbb{B}^{\star})+C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$
which is the desired inequality. ∎
D.1 Empirical Verification of the PL Inequality
An open question in the literature is the verification of the PL inequality (Equation 42) for general measures, such as empirical ones. To give further context, current results prove that it indeed holds for Gaussian measures, under specific spectral conditions chewi2020gradient. Therefore, we do an empirical validation on whether this bound holds in the empirical case. In a nutshell, to test Equation 42, we need to compute the PL ratio,
$\text{PL}_{\text{ratio}}(\hat{P}_{\tau})=\dfrac{\lVert\nabla_{W}\mathbb{B}(\hat{P}_{\tau})\rVert_{L_{2}(\hat{P}_{\tau})}}{\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}},$ (49)
and verify that it indeed remains strictly positive over the course of our optimization. Computing the PL ratio in Equation 49 requires knowledge of $\mathbb{B}^{\star}$, which is, in general, intractable. However, for location-scatter families, it is known in closed form, since $P^{\star}$ itself is known. Therefore, we verify it in the context of the Swiss-roll measure, where all quantities are known. We show our results in Figure 7, which shows that the ratio converges with iterations to approximately $0.56$.
Figure 7: Empirical test of the PL inequality. From left to right, we show: (i) the optimality gap, (ii) the gradient norm, and (iii) the PL ratio as specified in Equation 49.
D.2 PL Inequality on Location-Scatter Family
An open question in the literature is the verification of the PL inequality (Equation 42) for general measures. Current results establish that it holds for Gaussian measures under specific spectral conditions [chewi2020gradient]. In this section, we show that the PL inequality transfers from Gaussians to the broader class of location-scatter families, which includes our Swiss roll experiments (Section 4).
Location-Scatter Families. Following alvarez2016fixed, let $Q_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ be a measure with mean $\mu_{0}$ and covariance $\Sigma_{0}$. The location-scatter family generated by $Q_{0}$ is,
$\mathcal{L}(Q_{0}):=\left\{T_{\sharp}Q_{0}:T(x)=Ax+m,b\in\mathbb{R}^{d},A\in\mathcal{S}^{++}_{d}\right\}$, (50)
where $\mathcal{S}^{++}_{d}$ denotes the cone of $d\times d$ symmetric positive definite matrices. Each $Q\in\mathcal{L}(Q_{0})$ has mean $b$ and covariance $A$, but inherits the higher-order structure of $Q_{0}$.
Isometry to the Bures-Wasserstein Manifold. The key insight comes from alvarez2016fixed: for any $Q_{1},Q_{2}\in\mathcal{L}(Q_{0})$, the 2-Wasserstein distance equals
$\mathbb{W}_{2}^{2}(Q_{1},Q_{2})=\|\mu_{1}-\mu_{2}\|^{2}+\text{Tr}(\Sigma_{1}+\Sigma_{2}-2(\Sigma_{2}^{1/2}\Sigma_{1}\Sigma_{2}^{1/2})^{1/2}),$ (51)
This is precisely the Wasserstein distance on the Bures-Wasserstein manifold: $\mathcal{BW}=\{\mathcal{N}(\mu,\Sigma):\mu\in\mathbb{R}^{d},\Sigma\in\mathcal{S}_{d}^{++}\}$ [takatsu2011wasserstein]. Henceforth, we note the elements of $\mathcal{BW}$ by $G_{\mu,\Sigma}$. In that sense $G_{\mu_{1},\Sigma_{1}}$ and $G_{\mu_{2},\Sigma_{2}}$ denote the Gaussian measures matching the moments of $Q_{1},Q_{2}\in\mathcal{L}(Q_{0})$. As a consequence, the restriction of $\mathbb{W}_{2}$ to $\mathcal{L}(Q_{0})$ is isometric to the Bures-Wasserstein manifold. Moreover, the optimal transport map between elements of $\mathcal{L}(Q_{0})$ takes the same form as for Gaussians:
$T^{\star}_{Q_{1}\to Q_{2}}(x)=\mu_{2}+A(x-\mu_{1}),\quad A=\Sigma_{1}^{-1/2}\left(\Sigma_{1}^{1/2}\Sigma_{2}\Sigma_{1}^{1/2}\right)^{1/2}\Sigma_{1}^{-1/2}.$ (52)
Barycenter Functional Equivalence. Let $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}\subset\mathcal{L}(Q_{0})$ with $Q_{k}=T_{k}Q_{0}$, such that $\mu_{k}$ and $\Sigma_{k}$ are the mean and covariance of $Q_{k}$. Let $\mathcal{G}=\{G_{\mu_{k},\Sigma_{k}}\}_{k=1}^{K}$ be the corresponding Gaussian measures. Equation 51 establishes:
$\mathbb{B}(P|\mathcal{Q})=\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\text{, where }P\in\mathcal{L}(Q_{0})\text{, and }G_{\mu,\Sigma}\in\mathcal{BW},$ (53)
where $\mu,\Sigma$ are the mean and covariance of $P\in\mathcal{L}(Q_{0})$. Furthermore, by alvarez2016fixed, the barycenter $P^{\star}\in\mathcal{L}(Q_{0})$ corresponds to the Bures-Wasserstein barycenter: if $P^{\star}$ is the minimizer of $P\mapsto\mathbb{B}(P|\mathcal{Q})$ over $\mathcal{L}(Q_{0})$ with mean $\mu^{\star}$ and covariance $\Sigma^{\star}$, then $G_{\mu^{\star},\Sigma^{\star}}$ is the minimizer of $G_{\mu,\Sigma}\mapsto\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})$.
Wasserstein Gradient Structure. The Wasserstein gradient of $\mathbb{B}$ at $P$ with mean and covariance $\mu$ and $\Sigma$ is (cf. Equation 40)
$\nabla_{W}\mathbb{B}(P|\mathcal{Q})=2\sum_{k=1}^{K}\lambda_{k}\left(\mathrm{Id}-T^{\star}_{P\to Q_{k}}\right).$ (54)
Since each $T^{\star}_{P\to Q_{k}}$ is affine (Equation 52), the gradient $\nabla_{W}\mathbb{B}(P|\mathcal{Q})$ is itself an affine map:
$\nabla_{W}\mathbb{B}(P|\mathcal{Q})(x)$ $=2\sum_{k=1}^{K}\lambda_{k}(x-(A_{k}x+b_{k})),$
$=2\biggr(x-(\sum_{k=1}^{K}\lambda_{k}A_{k})(x)+\sum_{k=1}^{K}\lambda_{k}b_{k}\biggr),$
$=\underbrace{\biggr(2(\text{Id}-\sum_{k=1}^{K}\lambda_{k}A_{k})\biggr)}_{M}(x)+\underbrace{2\sum_{k=1}^{K}\lambda_{k}b_{k}}_{c},$
$=Mx+c,$ (55)
where $M=2(\text{Id}-\bar{A})$, $c=2(\bar{A}\mu-\bar{\mu})$, where $\bar{A}=\sum_{k=1}^{K}\lambda_{k}A_{k}$ and $\bar{\mu}=\sum_{k=1}^{K}\lambda_{k}\mu_{k}$. More importantly, by Equation (52), this is the same affine map as for the Gaussian case, resulting in,
$\nabla_{W}\mathbb{B}(P|\mathcal{Q})=\nabla_{W}\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\quad\text{as vector fields }\mathbb{R}^{d}\to\mathbb{R}^{d}.$ (56)
We now establish that the $L^{2}$-norms of these gradients are also equal, which is the key step for transferring the PL inequality.
Lemma D.1 ($L^{2}$-norm of Affine Maps).
Let $T(x)=Mx+c$ be an affine map with $M\in\mathbb{R}^{d\times d}$ and $c\in\mathbb{R}^{d}$. For any probability measure $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with mean $\mu_{P}$ and covariance $\Sigma_{P}$,
$\|T\|_{L^{2}(P)}^{2}:=\int_{\mathbb{R}^{d}}\|T(x)\|^{2}\,dP(x)=\text{Tr}(M^{T}M\Sigma_{P})+\|M\mu_{P}+c\|^{2}.$ (57)
In particular, $\|T\|_{L^{2}(P)}^{2}$ depends on $P$ only through its first two moments.
Proof.
Expanding the squared norm:
$\|T\|_{L^{2}(P)}^{2}$ $=\int\|Mx+c\|^{2}\,dP(x)$
$=\int\|M(x-\mu_{P})+(M\mu_{P}+c)\|^{2}\,dP(x)$
$=\int\|M(x-\mu_{P})\|^{2}\,dP(x)+2\langle M\mu_{P}+c,M\underbrace{\int(x-\mu_{P})\,dP(x)}_{=0}\rangle+\|M\mu_{P}+c\|^{2}$
$=\int(x-\mu_{P})^{T}M^{T}M(x-\mu_{P})\,dP(x)+\|M\mu_{P}+c\|^{2}$
$=\text{Tr}(M^{T}M\Sigma_{P})+\|M\mu_{P}+c\|^{2},$
where the last step uses $\mathbb{E}_{x\sim P}[(x-\mu_{P})(x-\mu_{P})^{\top}]=\Sigma_{P}$. ∎
Lemma D.2 (Gradient Norm Equality).
For any $P\in\mathcal{L}(Q_{0})$ and its corresponding Gaussian $G_{\mu,\Sigma}$:
$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P)}^{2}=\|\nabla_{W}\mathbb{B}(G|\mathcal{G})\|_{L^{2}(G_{\mu,\Sigma})}^{2}.$ (58)
Proof.
First, define $T_{\text{LS}}=\nabla_{W}\mathbb{B}(P|\mathcal{Q})$ and $T_{\text{Gaussian}}=\nabla_{W}\mathbb{B}(G_{m,\Sigma}|\mathcal{G})$. By construction, both of these mappings are affine, and by Equation 56 $T_{\text{LS}}=T_{\text{Gaussian}}$. From Lemma D.2, since $P$ and $G_{\mu,\Sigma}$ share the first and second order moments by construction, they are equal, which means,
$\lVert\nabla_{W}\mathbb{B}(P|\mathcal{Q})\rVert_{L_{2}(P)}^{2}=T_{\text{LS}}=T_{\text{Gaussian}}=\lVert\nabla_{W}\mathbb{B}(G|\mathcal{G})\rVert_{L_{2}(G_{\mu,\Sigma})}^{2},$
which is the desired result. ∎
We can now state and prove the following result.
Proposition D.2 (PL Inequality Transfer).
Let $\mathcal{L}(Q_{0})$ be a location-scatter family and $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}\subset\mathcal{L}(Q_{0})$. Let $\mathcal{G}=\{G_{\mu_{k},\Sigma_{k}}\}_{k=1}^{K}$ be the corresponding Gaussian with matching moments. If the Bures-Wasserstein barycenter functional $\mathbb{B}(\cdot|\mathcal{G})$ satisfies the PL inequality with constant $C_{\text{PL}}>0$, i.e.,
$\|\nabla_{W}\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\|_{L^{2}(G_{\mu,\Sigma})}^{2}\geq C_{\text{PL}}\left(\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})-\mathbb{B}(G_{\mu^{\star},\Sigma^{\star}}|\mathcal{G})\right),$ (59)
then $\mathbb{B}(\cdot|\mathcal{Q})$ satisfies the PL inequality on $\mathcal{L}(Q_{0})$ with the same constant,
$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P_{m,\Sigma})}^{2}\geq C_{\mathrm{PL}}\left(\mathbb{B}(P|\mathcal{Q})-\mathbb{B}(P^{\star}|\mathcal{Q})\right).$ (60)
Proof.
Starting from the PL inequality for Gaussians (Equation 59):
$\|\nabla_{W}\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\|_{L^{2}(G_{\mu,\Sigma})}^{2}\geq C_{\text{PL}}\left(\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})-\mathbb{B}(G_{\mu^{\star},\Sigma^{\star}}|\mathcal{G})\right).$
Applying Lemma D.2 to the left-hand side:
$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P)}^{2}\geq C_{\text{PL}}\left(\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})-\mathbb{B}(G_{\mu^{\star},\Sigma^{\star}}|\mathcal{G})\right).$
Applying Equation (53) and barycenter correspondence to the right-hand side:
$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P)}^{2}\geq C_{\mathrm{PL}}\left(\mathbb{B}(P|\mathcal{Q})-\mathbb{B}(P^{\star}|\mathcal{Q})\right).\qed$
Explicit PL Constant. By chewi2020gradient, the PL inequality holds for Gaussians under $\zeta$-regularity: if $\|\Sigma_{k}\|_{\mathrm{op}}\leq 1$ and $\det\Sigma_{k}\geq\zeta$ for all $k$, then $C_{\mathrm{PL}}=\zeta^{2}/4$. Then, Proposition D.2 extends this PL constant to location-scatter families satisfying the same spectral bounds.
Scope and Limitations. Proposition D.2 applies to location-scatter families, which are generated by affine transformations of a base measure. This includes elliptical distributions (e.g., multivariate $t$-distributions, uniform distributions on ellipsoids) and, importantly, the Swiss roll measures used in our experiments. This result does not directly apply to arbitrary measures.
Appendix E Domain Adaptation
E.1 Introduction
@@
Statistical learning theory vapnik1999overview deals with the problem of learning a function $h:\mathcal{X}\rightarrow\mathcal{Y}$, such that it minimizes the risk of making a mistake, that is,
@@
There is a fundamental assumption at play in the theory of generalization under the ERM framework. Indeed, both $\mathcal{R}_{P}$ and $\hat{\mathcal{R}}_{P}$ evaluate the risk under a fixed probability measure $P$. However, machine learning models are oftentimes confronted with data following different, but related probability measures quinonero2022dataset, a process known as distribution shift.
@@
E.2 Source-Only Training
A straightforward baseline in domain adaptation is the source-only training. Let $h=g\circ f$ be the composition of a feature extractor $g:\mathcal{X}\rightarrow\mathcal{U}$ and a feature classifier $f:\mathcal{U}\rightarrow\mathcal{Y}$. We can train $h$ end-to-end by minimizing,
@@ E.2 Source-Only Training
This training produces an embedding space $\mathcal{U}$ that is discriminative of the classes in the domain adaptation problem. Following previous work montesuma2021wasserstein, montesuma2023multi, montesuma2024distillation, we perform adaptation in this space.
E.3 Wasserstein Barycenter Transport
The motivation for using Wasserstein barycenters in multi-source domain adaptation comes from a theoretical result of redko2019optimal. Indeed, by [redko2019optimal, Theorem 2],
@@ E.3 Wasserstein Barycenter Transport
Plugging this back into inequality 64,
@@
4:  Compute $P^{\star}=\text{Bar}(\lambda,\mathcal{Q})$
5:  Compute $\gamma^{\star}$ between $P^{\star}$ and $Q_{T}$ (eq. 65)
6:  Transport $(x_{i}^{(P)},y_{i}^{(P)})$ (eq. 66)
@@
Wasserstein Barycenter Transport $(p=2)$. montesuma2021wasserstein uses a two-step procedure for finding $P$. First, one solves,